OpenStudy (anonymous):

Solving a problem about the eigenvalues of an operator and the probability of each outcome

OpenStudy (anonymous):

I saw this example here: http://quantummechanics.ucsd.edu/ph130a/130_notes/node261.html#example:Lxev I can't figure out why they set the eigenvalue to be \[\frac{ \hbar }{ \sqrt{2} }*b\]. Do they take into account that this is an angular momentum projection operator? How about solving the problem about any other operator?

OpenStudy (fwizbang):

In the example you sent, \[ L_x = \frac{\hbar}{\sqrt{2}} \] times the matrix. and the eigenvalue they are seeking is really $a$. They've moved the $\hbar$ factor to the other side and defined $b$ to simplify the algebra.